# Rare and exclusive few-body decays of the Higgs, Z, W bosons, and the top quark

### v3

@@ VI.1 Two-body top quark decays
The top quark decays to a W boson and a bottom quark, $\rm t\to W\,b$, with a branching fraction of nearly 100%, with the other tree-level decays $\rm t\to W\,s$ and $\rm t\to W\,d$ comparatively suppressed by factors of $10^{-3}$ and $10^{-4}$, respectively, as per the CKM element hierarchy $V_{\mathrm{tb}}\gg V_{\mathrm{ts}}>V_{\mathrm{td}}$ (Table 2). The FCNC top-quark decays to a gauge boson plus a light up-type quark (q = u or c; for charge conservation), $\rm t\to q\,Z$, $\rm t\to\gamma\,c$, and $\rm t\to g\,c$, occur only at the level of quantum loop corrections (Fig. 29), and are extremely suppressed in the SM. More precisely, they are 1-loop-, CKM-, and GIM-suppressed (with a GIM suppression factor [^13] Theoretically, the b-quark running mass must be evaluated at the top pole mass scale (at which the t quark decays) in the $\rm\overline{MS}$ scheme, otherwise rates are artificially enhanced. given by $f(m_{\mathrm{b}}^{2}/m_{\mathrm{W}}^{2})\approx 10^{-9}$), leading to tiny amplitudes. In many BSM models, however, the GIM suppression can be relaxed, and one-loop diagrams mediated by new bosons may also contribute, yielding effective couplings orders of magnitude larger, and correspondingly enhancing such ultrarare branching fractions [27, 28].
@@
Table 23: Compilation of rare two-body top quark decays. For each decay, we provide the predicted branching fraction(s) and the theoretical approach used to compute it, as well as the current experimental upper limit and that estimated for HL-LHC. The last two columns indicate whether the decay can be produced at FCC-ee/FCC-hh.
@@
$\rm t\;\to$ V $+$ q Branching fraction Framework 2024 HL-LHC FCC-ee FCC-hh
t\;$\to$ $\rm\gamma$ + $\rm c$  $(4.55\pm 0.23)$$\times~{}10^{-14}$ NLO SM [180] $<$ $4.0$×$\times~{}10^{-4}$ [181] $\lesssim 5.2\times~{}10^{-5}$ [62]
 $\left(4.6^{+2.0}_{-1.0}\right)$$\times~{}10^{-14}$ NLO SM [28]
$\rm u$  $(3.26\pm 0.34)$$\times~{}10^{-16}$ NLO SM [180] $<$ $8.9$×$\times~{}10^{-5}$ [181] $\lesssim 6.1\times~{}10^{-6}$ [62]
 $3.7$$\times~{}10^{-16}$ NLO SM [28]
$\rm g$ + $\rm c$  $(5.31\pm 0.27)$$\times~{}10^{-12}$ NLO SM [180] $<$ $4.1$×$\times~{}10^{-4}$ [182] $\lesssim 2.3\times~{}10^{-5}$ [62]
 $5.7$$\times~{}10^{-12}$ NLO SM [183]
 $\left(4.6^{+3.0}_{-1.0}\right)$$\times~{}10^{-12}$ NLO SM [28]
$\rm u$  $(3.81\pm 0.34)$$\times~{}10^{-14}$ NLO SM [180] $<$ $2.0$×$\times~{}10^{-5}$ [182] $\lesssim 2.7\times~{}10^{-6}$ [62]
 $3.7$$\times~{}10^{-14}$ NLO SM [28]
$\rm Z$ + $\rm c$  $1.0$$\times~{}10^{-14}$ NLO SM [28] $<$ $2.4$×$\times~{}10^{-4}$ [141] $\lesssim 2.3\times~{}10^{-5}$ [63]
$\rm u$  $8.0$$\times~{}10^{-17}$ NLO SM [28] $<$ $1.7$×$\times~{}10^{-4}$ [141] $\lesssim 7.3\times~{}10^{-6}$ [63]
$\rm H$ + $\rm c$  $\left(4.19^{+1.09}_{-0.86}\right)$$\times~{}10^{-15}$ NLO SM [184] $<$ $7.3$×$\times~{}10^{-4}$ [185] $\lesssim 8.5\times~{}10^{-5}$ [64]
 $3.0$$\times~{}10^{-15}$ NLO SM [28]
$\rm u$  $\left(3.66^{+1.15}_{-0.97}\right)$$\times~{}10^{-17}$ NLO SM [184] $<$ $1.9$×$\times~{}10^{-4}$ [185] $\lesssim 8.5\times~{}10^{-5}$ [64]
 $2.0$$\times~{}10^{-17}$ NLO SM [28]
@@ VI.2 Three-body top quark decays
Thanks to its large mass, the top quark has different rare decays kinematically accessible involving the presence of multiple heavy bosons in the final state [186, 27]. The possibility of 3-body radiative decays of the top quark $\rm t\to W\,b\,X$, where X can be a Z or a Higgs boson has been considered, e.g., in Refs. [187, 188, 189] (the cases $\rm X=\gamma,g$ are simply NLO real QED or QCD corrections to the dominant $\rm tW\,b$ decay, for which one needs also an energy threshold to avoid infrared/collinear divergences in the decay rates, and not considered here as they are not rare). The corresponding diagrams are shown in Fig. 31. Properly taking into account the finite width effects is key to compute any close-to-threshold 3-body decays with (partially) offshell particles [189, 183].
@@
Table 24 collects the branching fractions for rare three-body decays of the top quark, and Fig. 32 presents them in graphical form. No experimental search has been performed to date. By an amusing coincidence, the $\rm t\to Z\,W\,b$ decay is kinematically allowed because $m_{\mathrm{t}}\approx(m_{\mathrm{W}}+m_{\mathrm{b}})+m_{\mathrm{Z}}$ to within a few GeV, which can be satisfied within the $\mathcal{O}(2$ GeV) widths of the onshell electroweak bosons (Table 2). The corresponding branching fraction is $\mathcal{B}(\rm t\to Z\,W\,b)\approx 2\cdot 10^{-6}$ and, although no experimental limits exist yet from the LHC data, it could be potentially discovered at the HL-LHC, and for sure observed at FCC-ee and FCC-hh. The branching fraction of $\rm t\to H\,W\,b$, first evaluated in [190], has been recomputed here at NLO using MG5_aMC (via $\rm t\to H\,W^{*}(\mu\nu)b$, with $m_{\mathrm{b}}(m_{\mathrm{t}})=2.6$ GeV) with the parameters listed in Table 2, and found to be $\mathcal{B}=1.6\cdot 10^{-9}$. Other decays are highly GIM-suppressed (Fig. 31, bottom) and have much smaller rates. The $\rm t\to ZZq$ is below $\mathcal{O}(10^{-13})$, and is likely impossible to observe anywhere unless some BSM mechanism enhances it. The top quark can also decay into three up-type quarks, either through the same diagrams shown in Fig. 29 where the emitted boson further decays/splits into $\mathrm{u}\overline{\mathrm{u}}$ or $\mathrm{c}\overline{\mathrm{c}}$, or through a virtual W boson exchange. This process is dominated by the gluon splitting into two up-type quarks shown in the Fig. 31 (lower right panel), and has a decay rate of $\mathcal{O}(10^{-12})$ [191] commensurate with that of the “parent” $\rm t\to u\,g$ two-body decay (Table 23).
Table 24: Compilation of rare three-body top quark decays ($\rm u_{1,2}=u,c$ quarks). For each decay, we provide the predicted branching fraction(s) and the theoretical approach used to compute it. The last two columns indicate whether the decay can be produced at FCC-ee/FCC-hh.
@@
$\rm t\,\to$ X + Y + Z Branching fraction Framework 2024 HL-LHC FCC-ee FCC-hh
t\,$\to$ $\rm W^{+}$ + $\rm W^{-}$ + $\rm c$  $2.0$$\times~{}10^{-13}$ LO SM [192]
 $1.0$$\times~{}10^{-13}$ LO SM [27]
$\rm Z$ + $\rm W^{+}$ + $\rm b$  $1.8$$\times~{}10^{-6}$ LO SM [193]
 $2.0$$\times~{}10^{-6}$ LO SM [189]
 $\left(5.4^{+4.7}_{-2.0}\right)$$\times~{}10^{-7}$ LO SM [194]
$\rm H$ + $\rm W^{+}$ + $\rm b$  $1.8$$\times~{}10^{-9}$ LO SM [190]
 $1.6$$\times~{}10^{-9}$ NLO SM, MG5_aMC (this work)
$\rm Z$ + $\rm Z$ + $\rm c$ $<$ $1.0$$\times~{}10^{-13}$ LO SM [192]
$\rm u_{1}$ + $\rm u_{2}$ + $\rm\overline{u}_{2}$  $3.4$$\times~{}10^{-12}$ NLO SM [191]
@@ VI.3 Semiexclusive top quark decays into a quark plus a meson
There are also theoretical studies of semiexclusive top quark decays in which interactions among the decay quarks (either from the W decay, or combining the primary bottom quark with the W-decay quarks) lead to the formation of final states with one meson recoiling against a jet [26, 195, 196, 197]. Such decays can provide a new method to measure the top-quark mass via a two-body (jet-meson) invariant mass analysis, with different systematic uncertainties of those of the currently existing approaches [26]. Typical diagrams for the process are shown in Fig. 33 with the single meson recoiling against a $\rm q=u,c$ quark (right), or against a b quark (left).
@@
Branching fractions are listed in Table 25 for a bottom meson plus a c or u quark. The results of this table are shown in graphical form in Fig. 34. Theoretical branching fractions are in the $\mathcal{O}(10^{-5}$$10^{-12})$ range. No experimental study has been performed to date. The combined $\rm t\to B_{\mathrm{(s)}}+u/c$ decays have a relatively large rate of $\mathcal{B}\approx 10^{-4}$, that makes them worth an experimental search at the LHC. Dedicated studies exist that prove the feasibility to observe them via $\rm t\to b$-jet+\,+\,+c-jet at the HL-LHC [26]. Four channels are producible at FCC-ee, and all of them can be produced at FCC-hh.
Table 25: Compilation of semiexclusive decays of the top quark into a c or u quark plus a meson. For each decay, we provide the predicted branching fraction(s) and the theoretical approach used to compute it, as well as the current experimental upper limit and that estimated for HL-LHC. The last two columns indicate whether the decay can be produced at FCC-ee/FCC-hh.
@@
$\rm t\;\to$ M $+$ q Branching fraction Framework 2024 HL-LHC FCC-ee FCC-hh
t\;$\to$ $\rm\overline{B}^{0}$ + $\rm c$  $\left(2.1^{+2.1}_{-1.1}\right)$$\times~{}10^{-6}$ NRQCD+LDME [26]
$\rm u$  $\left(4.0^{+4.0}_{-2.0}\right)$$\times~{}10^{-5}$ NRQCD+LDME [26]
$\rm\overline{B}^{0}_{s}$ + $\rm c$  $\left(4.0^{+4.0}_{-2.0}\right)$$\times~{}10^{-5}$ NRQCD+LDME [26] $\lesssim 4.0\times~{}10^{-5}$ [26]
$\rm u$  $\left(2.1^{+2.1}_{-1.1}\right)$$\times~{}10^{-6}$ NRQCD+LDME [26]
$\rm\Upsilon(1S)$ + $\rm c$  $4.3$$\times~{}10^{-10}$ NRQCD+CSM [197]
@@
 $(6.4\pm 1.3)$$\times~{}10^{-10}$ NRQCD+COM [196]
[…]
$\rm\Upsilon(2S)$ + $\rm c$  $2.1$$\times~{}10^{-10}$ NRQCD+CSM [197]
[…]
$\rm\Upsilon(3S)$ + $\rm c$  $1.6$$\times~{}10^{-10}$ NRQCD+CSM [197]
[…]
$\rm\Upsilon(nS)$ + $\rm c$  $\left(1.9^{+0.2}_{-0.1}\right)$$\times~{}10^{-9}$ NRQCD+LDME [26]
$\rm u$  $\left(1.9^{+0.2}_{-0.1}\right)$$\times~{}10^{-11}$ NRQCD+LDME [26]
@@
The right diagram of Fig. 33 corresponds to the process of a top quark decaying through an offshell W boson with mass $m_{\mathrm{W^{*}}}\approx m_{\mathrm{M}}$ close to a mesonic resonance M that recoils against a b-quark jet. The amplitude for the process $\rm t\to b+M^{+}$ is similar to that for the well-known $\rm\tau^{\pm}\to\pi^{\pm}+\nu_{\tau}$ and $\rm\tau^{\pm}\to\rho^{\pm}+\nu_{\tau}$ processes [44]
@@
where $b_{\mathrm{M}}=0$ for pseudoscalar mesons ($\rm M^{\pm}=\pi^{\pm}$) and $b_{\mathrm{M}}=2$ for vector mesons ($\rm M^{\pm}=\rho^{\pm}$), and the factor $S_{\text{EW}}=1.0154$ includes the leading-logarithmic [198, 199] and non-logarithmic electroweak corrections [200]. For the particular $\rm t\to b+\pi^{+}/\rho^{+}$ decays, the relevant amplitudes read
@@
where $f_{\mathrm{M}}$ is the decay constant of the meson, and $|V_{\mathrm{qq^{\prime}}}|^{2}$ the relevant CKM matrix element, and where the last numerical equality is obtained with the parameters of Table 2. The decay $\mathcal{B}(\rm t\to\pi^{+}+b)$ was first estimated in Ref. [27], but their expression differs from Eq. (53) here by a factor of 1/9 [201]. We extend here the study of semiexclusive $\rm t\to meson+b$-jet decays to include the $\rm M^{\pm}=\rho^{\pm},\,K^{\pm},\,D^{\pm},\,D^{\pm}_{s},\,B^{\pm},\,B_{\mathrm{c}}^{\pm}$ channels (and their corresponding excited states) by employing the decay constants from the recent compilation for the light mesons [44], and from lattice calculations [70, 73] for the heavy ones (Table 2). The resulting $\rm t\to b+M$ branching fractions are tabulated in Table 26, and also shown in graphical form in Fig. 35. Estimated theoretical branching fractions are in the $\mathcal{O}(10^{-7}$$10^{-12})$ range. No experimental search has been performed to date at the LHC. Although no channel is producible in the clean FCC-ee environment, most channels will be accessible at FCC-hh.
Table 26: Compilation of semiexclusive decays of the top quark into a b quark plus a meson. For each decay, we provide the predicted branching fraction(s) and the theoretical approach used to compute it, or derived here using Eq. (53). The last two columns indicate whether the decay can be produced at FCC-ee/FCC-hh.
@@
$\rm t\;\to$ M $+$ b Branching fraction Framework 2024 HL-LHC FCC-ee FCC-hh
t\;$\to$ $\rm\pi^{\pm}$ + $\rm b$  $2.0$$\times~{}10^{-8}$ EFT+LCDA [27]
 $1.6$$\times~{}10^{-7}$ EFT+LCDA (this work)
$\rm\rho^{\pm}$ + $\rm b$  $4.3$$\times~{}10^{-7}$ EFT+LCDA (this work)
$\rm K^{\pm}$ + $\rm b$  $1.2$$\times~{}10^{-8}$ EFT+LCDA (this work)
$\rm K^{*\pm}$ + $\rm b$  $2.1$$\times~{}10^{-8}$ EFT+LCDA (this work)
$\rm D^{\pm}$ + $\rm b$  $2.3$$\times~{}10^{-8}$ EFT+LCDA (this work)
$\rm D^{*\pm}$ + $\rm b$  $3.7$$\times~{}10^{-8}$ EFT+LCDA (this work)
$\rm D_{s}^{\pm}$ + $\rm b$  $1.0$$\times~{}10^{-7}$ EFT+LCDA [27]
 $6.0$$\times~{}10^{-7}$ EFT+LCDA (this work)
$\rm D_{s}^{*\pm}$ + $\rm b$  $9.8$$\times~{}10^{-7}$ EFT+LCDA (this work)
$\rm B^{\pm}$ + $\rm b$  $5.3$$\times~{}10^{-12}$ EFT+LCDA (this work)
$\rm B^{*\pm}$ + $\rm b$  $4.7$$\times~{}10^{-12}$ EFT+LCDA (this work)
$\rm B_{c}^{\pm}$ + $\rm b$  $3.1$$\times~{}10^{-9}$ EFT+LCDA (this work)
$\rm B_{c}^{*\pm}$ + $\rm b$  $3.1$$\times~{}10^{-9}$ EFT+LCDA (this work)

### v4

@@ VI.1 Two-body top quark decays
The top quark decays to a W boson and a bottom quark, $\rm t\to W\,b$, with a branching fraction of nearly 100%, with the other tree-level decays $\rm t\to W\,s$ and $\rm t\to W\,d$ comparatively suppressed by factors of $10^{-3}$ and $10^{-4}$, respectively, as per the CKM element hierarchy $V_{\mathrm{tb}}\gg V_{\mathrm{ts}}>V_{\mathrm{td}}$ (Table 2). The FCNC top-quark decays to a gauge boson plus a light up-type quark (q = u or c; for charge conservation), $\rm t\to q\,Z$, $\rm t\to\gamma\,c$, and $\rm t\to g\,c$, occur only at the level of quantum loop corrections (Fig. 29), and are extremely suppressed in the SM. More precisely, they are 1-loop-, CKM-, and GIM-suppressed (with a GIM suppression factor [^15] Theoretically, the b-quark running mass must be evaluated at the top pole mass scale (at which the t quark decays) in the $\rm\overline{MS}$ scheme, otherwise rates are artificially enhanced. given by $f(m_{\mathrm{b}}^{2}/m_{\mathrm{W}}^{2})\approx 10^{-9}$), leading to tiny amplitudes. In many BSM models, however, the GIM suppression can be relaxed, and one-loop diagrams mediated by new bosons may also contribute, yielding effective couplings orders of magnitude larger, and correspondingly enhancing such ultrarare branching fractions [Beneke:2000hk, Aguilar-Saavedra:2004mfd].
@@
Table 23: Compilation of rare two-body top quark decays. For each decay, we provide the predicted branching fraction(s) and the theoretical approach used to compute it, as well as the current experimental upper limit and that estimated for HL-LHC. The last two columns indicate whether the decay can be produced at FCC-ee/FCC-$h$.
@@
$\rm t\;\to$ V ++ q Branching fraction Framework 2024 HL-LHC FCC-$e$ FCC-$h$
[…]
$\left(4.6^{+2.0}_{-1.0}\right)$ $\times~10^{-14}$ loop SM [Aguilar-Saavedra:2004mfd]
[…]
$3.7$ $\times~10^{-16}$ loop SM [Aguilar-Saavedra:2004mfd]
[…]
$5.7$ $\times~10^{-12}$ loop SM [Eilam:2006uh]
$\left(4.6^{+3.0}_{-1.0}\right)$ $\times~10^{-12}$ loop SM [Aguilar-Saavedra:2004mfd]
[…]
$3.7$ $\times~10^{-14}$ loop SM [Aguilar-Saavedra:2004mfd]
[…]
$3.0$ $\times~10^{-15}$ loop SM [Aguilar-Saavedra:2004mfd]
[…]
$2.0$ $\times~10^{-17}$ loop SM [Aguilar-Saavedra:2004mfd]
@@ VI.2 Three-body top quark decays
Thanks to its large mass, the top quark has different rare decays kinematically accessible involving the presence of multiple heavy bosons in the final state [Jenkins:1996zd, Beneke:2000hk]. The possibility of 3-body radiative decays of the top quark $\rm t\to W\,b\,X$, where X can be a Z or a Higgs boson has been considered, e.g., in Refs. [Decker:1992wz, Mahlon:1994us, Altarelli:2000nt] (the cases $\rm X=\gamma,g$ are simply NLO real QED or QCD corrections to the dominant $\rm tW\,b$ decay, for which one needs also an energy threshold to avoid infrared/collinear divergences in the decay rates, and not considered here as they are not rare). The corresponding diagrams are shown in Fig. 31. Properly taking into account the finite width effects is key to compute any close-to-threshold 3-body decays with (partially) offshell particles [Altarelli:2000nt, Eilam:2006uh].
@@
Table 24 collects the branching fractions for rare three-body decays of the top quark, and Fig. 32 presents them in graphical form. No experimental search has been performed to date. By an amusing coincidence, the $\rm t\to Z\,W\,b$ decay is kinematically allowed because $m_{\mathrm{t}}\approx(m_{\mathrm{W}}+m_{\mathrm{b}})+m_{\mathrm{Z}}$ to within a few GeV, which can be satisfied within the $\mathcal{O}(2$ GeV) widths of the onshell electroweak bosons (Table 2). The corresponding branching fraction is $\mathcal{B}(\rm t\to Z\,W\,b)\approx 2\cdot 10^{-6}$ and, although no experimental limits exist yet from the LHC data, it could be potentially discovered at the HL-LHC, and for sure observed at FCC-ee and FCC-hh. The branching fraction of $\rm t\to H\,W\,b$, first evaluated in [Han:2013sea], has been recomputed here at NLO using MG5_aMC (via $\rm t\to H\,W^{*}(\mu\nu)b$, with $m_{\mathrm{b}}(m_{\mathrm{t}})=2.6$ GeV) with the parameters listed in Table 2, and found to be $\mathcal{B}=1.6\cdot 10^{-9}$. Other decays are highly GIM-suppressed (Fig. 31, bottom) and have much smaller rates. The $\rm t\to ZZq$ is below $\mathcal{O}(10^{-13})$, and is likely impossible to observe anywhere unless some BSM mechanism enhances it. The top quark can also decay into three up-type quarks, either through the same diagrams shown in Fig. 29 where the emitted boson further decays/splits into $\mathrm{u}\overline{\mathrm{u}}$ or $\mathrm{c}\overline{\mathrm{c}}$, or through a virtual W boson exchange. This process is dominated by the gluon splitting into two up-type quarks shown in the Fig. 31 (lower right panel), and has a decay rate of $\mathcal{O}(10^{-12})$ [Cordero-Cid:2004pco] commensurate with that of the “parent” $\rm t\to u\,g$ two-body decay (Table 23).
Table 24: Compilation of rare three-body top quark decays ($\rm u_{1,2}=u,c$ quarks). For each decay, we provide the predicted branching fraction(s) and the theoretical approach used to compute it. The last two columns indicate whether the decay can be produced at FCC-ee/FCC-$h$.
@@
$\rm t\,\to$ X + Y + Z Branching fraction Framework 2024 HL-LHC FCC-$e$ FCC-$h$
[…]
$1.0$ $\times~10^{-13}$ LO SM [Beneke:2000hk]
$\rm Z$ + $\rm W^{+}$ + $\rm b$ $1.8$ $\times~10^{-6}$ LO SM [Papaefstathiou:2017xuv]
$2.0$ $\times~10^{-6}$ LO SM [Altarelli:2000nt]
$\left(5.4^{+4.7}_{-2.0}\right)$ $\times~10^{-7}$ LO SM [Mahlon:1998fr]
$\rm H$ + $\rm W^{+}$ + $\rm b$ $1.8$ $\times~10^{-9}$ LO SM [Han:2013sea]
$1.6$ $\times~10^{-9}$ loop SM, MG5_aMC (this work)
$\rm Z$ + $\rm Z$ + $\rm c$ << $1.0$ $\times~10^{-13}$ LO SM [Bar-Shalom:2005ldb]
$\rm u_{1}$ + $\rm u_{2}$ + $\rm\overline{u}_{2}$ $3.4$ $\times~10^{-12}$ loop SM [Cordero-Cid:2004pco]
@@ VI.3 Semiexclusive top quark decays into a quark plus a meson
There are also theoretical studies of semiexclusive top quark decays in which interactions among the decay quarks (either from the W decay, or combining the primary bottom quark with the W-decay quarks) lead to the formation of final states with one meson recoiling against a jet [dEnterria:2020ygk, Handoko:1998tz, Handoko:1999iu, Slabospitskii:2021rns]. Such decays can provide a new method to measure the top-quark mass via a two-body (jet-meson) invariant mass analysis, with different systematic uncertainties of those of the currently existing approaches [dEnterria:2020ygk]. Typical diagrams for the process are shown in Fig. 33 with the single meson recoiling against a $\rm q=u,c$ quark (right), or against a b quark (left).
@@
Branching fractions are listed in Table 25 for a bottom meson plus a c or u quark. The results of this table are shown in graphical form in Fig. 34. Theoretical branching fractions are in the $\mathcal{O}(10^{-5}$$10^{-12})$ range. No experimental study has been performed to date. The combined $\rm t\to B_{\mathrm{(s)}}+u/c$ decays have a relatively large rate of $\mathcal{B}\approx 10^{-4}$, that makes them worth an experimental search at the LHC. Dedicated studies exist that prove the feasibility to observe them via $\rm t\to b$-jet+\,+\,c-jet at the HL-LHC [dEnterria:2020ygk]. Four channels are producible at FCC-ee, and all of them can be produced at FCC-hh.
Table 25: Compilation of semiexclusive decays of the top quark into a c or u quark plus a meson. For each decay, we provide the predicted branching fraction(s) and the theoretical approach used to compute it, as well as the current experimental upper limit and that estimated for HL-LHC. The last two columns indicate whether the decay can be produced at FCC-ee/FCC-$h$.
@@
$\rm t\;\to$ M ++ q Branching fraction Framework 2024 HL-LHC FCC-$e$ FCC-$h$
@@
$\rm u$ $\left(4.0^{+4.0}_{-2.0}\right)$ $\times~10^{-5}$ NRQCD+LDME [dEnterria:2020ygk]
$\rm\overline{B}^{0}_{s}$ + $\rm c$ $\left(4.0^{+4.0}_{-2.0}\right)$ $\times~10^{-5}$ NRQCD+LDME [dEnterria:2020ygk] $\lesssim 4.0\times~10^{-5}$ [dEnterria:2020ygk]
$\rm u$ $\left(2.1^{+2.1}_{-1.1}\right)$ $\times~10^{-6}$ NRQCD+LDME [dEnterria:2020ygk]
$\rm\Upsilon(1S)$ + $\rm c$ $4.3$ $\times~10^{-10}$ NRQCD+CSM [Slabospitskii:2021rns]
$(1.0\text{ -- }1.5)$ $\times~10^{-9}$ NRQCD+LDME [dEnterria:2020ygk]
$(6.4\pm 1.3)$ $\times~10^{-10}$ NRQCD+COM [Handoko:1999iu]
$\rm u$ $(1.0\text{ -- }1.5)$ $\times~10^{-11}$ NRQCD+LDME [dEnterria:2020ygk]
$\rm\Upsilon(2S)$ + $\rm c$ $2.1$ $\times~10^{-10}$ NRQCD+CSM [Slabospitskii:2021rns]
$(1.7\text{ -- }5.3)$ $\times~10^{-10}$ NRQCD+LDME [dEnterria:2020ygk]
$\rm u$ $(1.7\text{ -- }5.3)$ $\times~10^{-12}$ NRQCD+LDME [dEnterria:2020ygk]
$\rm\Upsilon(3S)$ + $\rm c$ $1.6$ $\times~10^{-10}$ NRQCD+CSM [Slabospitskii:2021rns]
$(2.7\text{ -- }3.8)$ $\times~10^{-10}$ NRQCD+LDME [dEnterria:2020ygk]
$\rm u$ $(2.7\text{ -- }3.8)$ $\times~10^{-12}$ NRQCD+LDME [dEnterria:2020ygk]
$\rm\Upsilon(nS)$ + $\rm c$ $\left(1.9^{+0.2}_{-0.1}\right)$ $\times~10^{-9}$ NRQCD+LDME [dEnterria:2020ygk]
$\rm u$ $\left(1.9^{+0.2}_{-0.1}\right)$ $\times~10^{-11}$ NRQCD+LDME [dEnterria:2020ygk]
@@
The right diagram of Fig. 33 corresponds to the process of a top quark decaying through an offshell W boson with mass $m_{\mathrm{W^{*}}}\approx m_{\mathrm{M}}$ close to a mesonic resonance M that recoils against a b-quark jet. The amplitude for the process $\rm t\to b+M^{+}$ is similar to that for the well-known $\rm\tau^{\pm}\to\pi^{\pm}+\nu_{\tau}$ and $\rm\tau^{\pm}\to\rho^{\pm}+\nu_{\tau}$ decays [Grossman:2015cak]
@@
where $b_{\mathrm{M}}=0$ for pseudoscalar mesons ($\rm M^{\pm}=\pi^{\pm}$) and $b_{\mathrm{M}}=2$ for vector mesons ($\rm M^{\pm}=\rho^{\pm}$), and the factor $S_{\text{EW}}=1.0154$ includes the leading-logarithmic [Sirlin:1977sv, Marciano:1988vm] and non-logarithmic electroweak corrections [Braaten:1990ef]. For the particular $\rm t\to b+\pi^{+}/\rho^{+}$ decays, the relevant amplitudes read
@@
where $f_{\mathrm{M}}$ is the decay constant of the meson, and $|V_{\mathrm{qq^{\prime}}}|^{2}$ the relevant CKM matrix element, and where the last numerical equality is obtained with the parameters of Table 2. The decay $\mathcal{B}(\rm t\to\pi^{+}+b)$ was first estimated in Ref. [Beneke:2000hk], but their expression differs from Eq. (53) here by a factor of 1/9 [Long-ShunLu2024]. We extend here the study of semiexclusive $\rm t\to meson+b$-jet decays to include the $\rm M^{\pm}=\rho^{\pm},\,K^{\pm},\,D^{\pm},\,D^{\pm}_{s},\,B^{\pm},\,B_{\mathrm{c}}^{\pm}$ channels (and their corresponding excited states) by employing the decay constants from the recent compilation for the light mesons [Grossman:2015cak], and from lattice calculations [Becirevic:2012ti, Colquhoun:2015oha] for the heavy ones (Table 2). The resulting $\rm t\to b+M$ branching fractions are tabulated in Table 26, and also shown in graphical form in Fig. 35. Estimated theoretical branching fractions are in the $\mathcal{O}(10^{-7}$$10^{-12})$ range. No experimental search has been performed to date at the LHC. About four channels are producible in the clean FCC-ee environment, and all of them will be accessible at FCC-hh.
Table 26: Compilation of semiexclusive decays of the top quark into a b quark plus a meson. For each decay, we provide the predicted branching fraction(s) and the theoretical approach used to compute it, or derived here using Eq. (53). The last two columns indicate whether the decay can be produced at FCC-ee/FCC-$h$.
@@
$\rm t\;\to$ M ++ b Branching fraction Framework 2024 HL-LHC FCC-$e$ FCC-$h$
[…]
$2.0$ $\times~10^{-8}$ EFT+LCDA [Beneke:2000hk]
$\rm\rho^{\pm}$ + $\rm b$ $4.3$ $\times~10^{-7}$ Eq. (53) (this work)
$\rm K^{\pm}$ + $\rm b$ $1.2$ $\times~10^{-8}$ Eq. (53) (this work)
$\rm K^{*\pm}$ + $\rm b$ $2.1$ $\times~10^{-8}$ Eq. (53) (this work)
$\rm D^{\pm}$ + $\rm b$ $2.3$ $\times~10^{-8}$ Eq. (53) (this work)
$\rm D^{*\pm}$ + $\rm b$ $3.7$ $\times~10^{-8}$ Eq. (53) (this work)
$\rm D_{s}^{\pm}$ + $\rm b$ $6.0$ $\times~10^{-7}$ Eq. (53) (this work)
$1.0$ $\times~10^{-7}$ EFT+LCDA [Beneke:2000hk]
$\rm D_{s}^{*\pm}$ + $\rm b$ $9.8$ $\times~10^{-7}$ Eq. (53) (this work)
$\rm B^{\pm}$ + $\rm b$ $5.3$ $\times~10^{-12}$ Eq. (53) (this work)
$\rm B^{*\pm}$ + $\rm b$ $4.7$ $\times~10^{-12}$ Eq. (53) (this work)
$\rm B_{c}^{\pm}$ + $\rm b$ $3.1$ $\times~10^{-9}$ Eq. (53) (this work)
$\rm B_{c}^{*\pm}$ + $\rm b$ $3.1$ $\times~10^{-9}$ Eq. (53) (this work)